OpenStudy (anonymous):

Let f(x) = 4x2 + x + 1 and g(x) = x2 – 2. Find g(f(x)). Show each step of your work.

OpenStudy (decentnabeel):

g(f(x))=(4x^2 + x + 1)^2-2

OpenStudy (decentnabeel):

g(x)=x^2-1 put x=f(x)=4x^2 + x + 1 g(f(x))=(4x^2 + x + 1)^2-2

OpenStudy (decentnabeel):

are you understand @mollzlask

OpenStudy (anonymous):

yess! thank youuuu

OpenStudy (decentnabeel):

no problem :)